OpenStudy (anonymous):

Describe how the graph of y= x^2 can be transformed to the graph of the given equation. y = (x-4)^2-8

OpenStudy (anonymous):

Shift the graph of y = x^2 right 4 units and then up 8 units. Shift the graph of y = x^2 right 4 units and then down 8 units. i think it's this one Shift the graph of y = x^2 left 4 units and then down 8 units. Shift the graph of y = x^2 down 4 units and then left 8 units

OpenStudy (primeralph):

What do you think?

OpenStudy (anonymous):

2nd

OpenStudy (anonymous):

third

OpenStudy (anonymous):

two negative signs

OpenStudy (primeralph):

@yahya90 That's not allowed.

OpenStudy (jdoe0001):

y = A( Bx + C ) + D C = horizontal shift, C > 0, to the left, C < 0, to the right D = vertical shift, D > 0, up, D<0, down

OpenStudy (anonymous):

@primeralph so it's the second one !

OpenStudy (dumbcow):

@lornbeach haha basically everyone is confirming you are correct :)

OpenStudy (anonymous):

Thank youu (: